OpenStudy (anonymous):

Functions :) Homework problem. f(x) = 2x + 3 and g(x) = 9x - (1/2). Solve for f(14). I know how to solve it (the output of g(x) = the input of f(x)) but I forgot how to start... Is it by substituting 14 for x in g(x)? Can't remember.

OpenStudy (campbell_st):

well the question seems to ask for f(14) so substitute x = 14 into f(x) = 2x + 3 hope that makes sense

OpenStudy (anonymous):

So i'm thinking it's going to start like this? \[f(g(x)) = f(g(14))\] \[f(x()) = f(9(14) - (1/2))\] And that's how you start it? i can solve from there.

OpenStudy (campbell_st):

or is the question 1. g(f(14)) ? or 2. f(g(14)).?

OpenStudy (anonymous):

Well my teacher wrote it by hand so he might have meant find g(14)?

OpenStudy (anonymous):

If he wanted me to find f(14) it wouldn't make sense to give me the g(x) equation so i think he meant to say find g(14).

OpenStudy (campbell_st):

that's correct... but then it could have been included as a distraction... my advice... find g(14) then substitute the value into f(x) and also find f(14) and then substitute it into g(x) that covers most options

OpenStudy (anonymous):

Good plan. But back to the original question -- I forgot how to start out, so is my first statement correct? (as in f(g(x)) = f(9(14) - (1/2))

OpenStudy (campbell_st):

that's correct...

OpenStudy (anonymous):

Ok then, so... \[f(125.5)\] \[f(g(x)) = 2(125.5) + 3\] \[f(g(x)0 = 254\]

OpenStudy (anonymous):

There's an abstract zero in there o.O didn't mean to put that. Thanks camp!

OpenStudy (campbell_st):

that's great well done